# Optimal approximation of infinite-dimensional holomorphic functions II: recovery from i.i.d. pointwise samples

### v2

@@ 3 $m$-widths of $(\text{{{b}}},1)$-holomorphic functions: the results of [6]
Let $\mathcal{K}$ be a set of (holomorphic) functions. To understand how well we can approximate functions in $\mathcal{K}$ from finitely many samples, we use the theory of $m$-widths (see, e.g., [31]). Recall that $\mathcal{V}$ is a Hilbert space, $\varrho$ is a tensor-product probability measure on $\mathcal{U}$ and consider the Lebesgue–Bochner space $\mathcal{X}=L^{2}_{\varrho}(\mathcal{U};\mathcal{V})$. Let $\mathcal{Y}=C(\mathcal{U};\mathcal{V})$ be the normed vector space of continuous functions from $\mathcal{U}$ to $\mathcal{V}$ with respect to the uniform norm. Then, following [6], we define the (adaptive) $m$-width
@@
Now consider $(\bm{b},\varepsilon)$-holomorphic functions. In [6], we made the distinction between the known anisotropy (known $\bm{b}$) and unknown anisotropy (unknown $\bm{b}$) cases. Specifically, in the known anisotropy case, we let $\mathcal{K}=\mathcal{H}(\bm{b})$ in (3.1) for some fixed $\bm{b}$, so that the reconstruction map $\mathcal{T}$ can depend on $\bm{b}$. Conversely, in the case of unknown anisotropy we let $\mathcal{K}=\mathcal{H}(p)$ or $\mathcal{K}=\mathcal{H}(p,\mathsf{M})$ in (3.1), in which case the reconstruction map is prohibited from depending on $\bm{b}$.
@@
On the other hand, in the unknown anisotropy case, we write
@@
Note that the main results from [6] only considered the case where the probability measure $\varrho$ in (2.1) is a tensor product of the same measure, i.e., $\varrho_{1}=\varrho_{2}=\cdots$. However, they can be easily generalized subject to the following assumption (see Remark 3). In this assumption, we denote the standard deviation of $\varrho_{i}$ as
$\sigma_{i}=\sqrt{\int^{1}_{-1}(y-\tau_{i})^{2}\,\mathrm{d}\varrho_{i}(y)},\quad\text{where }\tau_{i}=\int^{1}_{-1}y\,\mathrm{d}\varrho_{i}(y).$ (3.5)
@@
    The probability measure $\varrho$ on $\mathcal{U}$ has the form (2.1), where the standard deviations (3.5) of the constituent measures $\varrho_{i}$ satisfy $\inf_{i}\sigma_{i}>0$.
@@
The bound (3.6) pertains to the known anisotropy case. It shows that no sampling-recovery pair can achieve a rate better than $m^{1/2-1/p}$ for $\bm{b}\in\ell^{p}(\mathbb{N})$ or $\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N})$ when the reconstruction map is allowed to depend on $\bm{b}$. For the unknown anisotropy case, (3.7) shows that it is impossible to approximate functions in $\mathcal{H}(p)$ from finitely-many samples. This arises because functions in $\mathcal{H}(p)$ can have important variables in infinitely many different configurations [6]. The final result (3.8) asserts that when restricting to the class $\mathcal{H}(p,\mathsf{M})$ one can achieve a rate of at best $m^{1/2-1/p}$.
We note in passing that [6, Thms. 4.6 & 4.7] also establish upper bounds for several of the $m$-widths considered above in the case where $\varrho$ is the uniform measure. However, these bounds do not consider i.i.d. pointwise samples and are therefore not relevant to the present paper. In the next section, we establish counterpart upper bounds based on using such samples. In particular, we show that the rates $m^{1/2-1/p}$ can be ‘nearly’ achieved from i.i.d. pointwise samples drawn from general tensor-product Jacobi measures, using suitable recovery maps based on the corresponding tensor Jacobi polynomials.
@@
    The extension of the results of [6] to the case where $\varrho$ satisfies Assumption 3 involves a straightforward modification of several proofs. First, one defines the orthonormal (in $L^{2}_{\varrho}(\mathcal{U})$) functions $\psi_{i}(\bm{y})=\frac{y_{i}-\tau_{i}}{\sigma_{i}}$ and then observes that [6, Lem. 5.2] remains valid with $\sigma=\inf_{i}\sigma_{i}$ and $\tau=\sup_{i}\tau_{i}$. Observe that $|\tau|\leq 1$ since each $\varrho_{i}$ is a probability measure supported on $[-1,1]$. Lemmas 5.3 and 5.4 of [6] then extend without further modifications and with these values of $\sigma$ and $\tau$. Finally, the same applies to Theorems 4.4 and 4.5 of [6], on which Theorem 3.2 is based.
@@
    In general, $\overline{\theta_{m}}(p)\neq\overline{\theta_{m}}(p,\mathsf{M})$. However, $\overline{\theta_{m}}(p)=\overline{\theta_{m}}(p,\mathsf{M})$ whenever the measure $\varrho$ (2.1) is a tensor product of the same measure, i.e., $\varrho_{1}=\varrho_{2}=\cdots$. This follows from two observations. First, for such a measure, we have $\theta_{m}(\bm{b})=\theta_{m}(\pi(\bm{b}))$ for any permutation $\pi$. Second, if $\bm{b}\in\ell^{p}(\mathbb{N})$ with ${\|\bm{b}\|}_{p}\leq 1$, then there is a permutation $\pi$ for which $\pi(\bm{b})$ is nonincreasing, and therefore $\pi(\bm{b})\in\ell^{p}_{\mathsf{M}}(\mathbb{N})$ with ${\|\pi(\bm{b})\|}_{p,\mathsf{M}}={\|\bm{b}\|}_{p}$. See Appendix B for additional discussion.
@@ 4.1 Jacobi measures
As noted in §1.1, most prior results for approximation of holomorphic functions from i.i.d. pointwise samples consider either the uniform or tensor-product Chebyshev/arcsine probability measure. Here, we consider the more general setting of tensor-product Jacobi measures. For $\alpha,\beta>-1$ the Jacobi measure on $[-1,1]$ is given by
@@
It is a short argument to show that Assumption 4.1 implies Assumption 3 for the measure (4.3). See Proposition A.1. Therefore, the lower bounds of Theorem 3.2 apply in this setting.
4.2 Jacobi polynomials
  • Remark 4.2
    for some polynomial $p$ that is independent of $j$ and depends on $\bm{\alpha}$ and $\bm{\beta}$ only. See [7, Rmk. 3.5] for a similar assumption in a related context. In other words, Assumption 4.1 ensures that the $L^{\infty}$-norms of the constituent Jacobi polynomials do not blow up too fast. This is a key ingredient in our proofs, which involve showing various weighted best approximation error bounds for Jacobi polynomial expansions of $(\bm{b},1)$-holomorphic functions (see Theorems 5.1 and 5.2), where the weights are based on these norms (see (5.8)).
@@ 4.3 Upper bounds in expectation
Let $\bm{\alpha},\bm{\beta}$ satisfy Assumption 4.1 with constant $\tau>0$ and $\varrho=\varrho_{\bm{\alpha},\bm{\beta}}$ be the tensor-product Jacobi probability measure as in (4.3). Let $\zeta>0$, $m\geq 3$ and consider sample points $\bm{y}_{1},\ldots,\bm{y}_{m}\sim_{\mathrm{i.i.d.}}\varrho$. Then there is a reconstruction map $\mathcal{T}:(\mathcal{U}\times\mathcal{V})^{m}\rightarrow L^{2}_{\varrho}(\mathcal{U};\mathcal{V})$ depending on $\zeta$, $\bm{\alpha}$ and $\bm{\beta}$ only such that
$\theta_{m}(\bm{b})\leq\mathbb{E}\left(\sup_{f\in\mathcal{H}(\bm{b})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\right)\leq{C(\bm{b},p,\zeta,{\tau})}\cdot\left(\frac{m}{\log^{4+\zeta}(m)}\right)^{\frac{1}{2}-\frac{1}{p}},$
@@
$\theta_{m}(p,\mathsf{M})\leq\mathbb{E}\left(\sup_{f\in\mathcal{H}(p,\mathsf{M})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\right)\leq{C(p,t,\zeta,{\tau})}\cdot m^{\frac{1}{2}-\frac{1}{t}},$
@@
Note that, in both Theorems 4.3 and 4.4, the leftmost inequalities (namely, $\theta_{m}(\bm{b})\leq\ldots$ and so forth) follow immediately from the definition of the relevant quantity, i.e., (3.2)–(3.4).
@@ 4.4 Discussion
We now discuss some important aspects of Theorems 4.34.4 and their consequences for the approximation theory of $(\bm{b},1)$-holomorphic functions from finite samples.
Theorem 3.2 provides lower bounds for the various $m$-widths, and establishes that the rate $m^{1/2-1/p}$ is the best possible. These lower bounds are valid for Banach-valued functions, whereas our main results hold only when $\mathcal{V}$ is a Hilbert space. This gap was previously encountered in [3]. Here, (nonuniform) upper bounds for Legendre and Chebyshev polynomial approximations were shown for Banach-valued functions. However, the algebraic rates of convergence shown therein are slower than those of the Hilbert-valued case. A similar extension of Theorems 4.3 and 4.4 to Banach spaces would also result in similar reductions of the rates. As of now, it is unknown how to attain upper bounds for the $m$-width for Banach-valued functions that match the lower bounds.
In the second parts of Theorems 4.3 and 4.4, we trade the (poly)logarithmic factor from the first part for an (arbitrarily small) reduction in the algebraic rate in order to obtain an upper bound for the corresponding $m$-width. This is done because it is not possible, with our current proof techniques, to bound the constant $C(\bm{b}_{\pi},p,\tau)$ (respectively, $C(\bm{b},p,\zeta,\tau)$) uniformly over sequences $\bm{b}\in\ell^{p}(\mathbb{N})$ with ${\|\bm{b}\|}_{p}\leq 1$ (respectively, $\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N})$ with ${\|\bm{b}\|}_{p,\mathsf{M}}\leq 1$). Instead, we use an idea from [6] which allows us to bound $\sup_{{\|\bm{b}\|}_{p}\leq 1}{C(\bm{b}_{\pi},t,\tau)}<\infty$ for $t\in(p,1)$ and likewise for $C(\bm{b},t,\zeta,{\tau})$. See the proofs of Theorems 4.3 and 4.4 in §6 for further details.
@@
As noted previously, the unknown anisotropy case studied in Theorem 4.4 requires a different approach to the least-squares approximation used in Theorem 4.3. Due to the intrinsic unknown nature of the entries of $\bm{b}$, we use techniques based on compressed sensing to establish theoretical guarantees on the recovery of certain polynomial coefficients in a large (but finite) index set. A key step involves using well-established results that assert the (weighted) Restricted Isometry Property (RIP) for the relevant measurement matrix. The need to assert the RIP leads to the $\log^{4}(m)$ factor in the first bound in Theorem 4.4. Unfortunately, it is well-known that improving the log factors in RIP-type estimates is a challenging open problem. See [10] for further discussion. The additional $\log^{\zeta}(m)$ term in the bound stems from the desire to have a reconstruction map $\mathcal{T}$ that is valid for all $10<p<$ 1. We remark in passing that $\log^{\zeta}(m)$ can be replaced by $g(m)$ for any (e.g., arbitrarily-slowly) increasing function $g:\mathbb{N}\rightarrow[1,\infty)$ with $g(m)\rightarrow+\infty$ as $m\rightarrow+\infty$, in which case the constant depends on $g$ instead of $\zeta$. See Remark 6.4 for further discussion on this term.
While Theorem 4.4 relies on compressed sensing, it is also possible to derive a weaker version of it by using the least-squares approximation from Theorem 4.3. Notice that any $\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N})$, $\bm{b}\geq\bm{0}$, ${\|\bm{b}\|}_{p,\mathsf{M}}\leq 1$ satisfies $\bm{b}\leq\bm{b}_{p}$, where $\bm{b}_{p}=(i^{-1/p})_{i\in\mathbb{N}}$. It follows from (2.4) that $\mathcal{R}(\bm{b}_{p},\varepsilon)\subseteq\mathcal{R}(\bm{b},\varepsilon)$, which implies that $\mathcal{H}(\bm{b})\subseteq\mathcal{H}(\bm{b}_{p})$ and, in turn, that $\theta_{m}(\bm{b})\leq\theta_{m}(\bm{b}_{p})$. Now consider the reconstruction map $\mathcal{T}=\mathcal{T}_{p}$ from Theorem 4.3 based on $\bm{b}_{p}$. Since $\bm{b}_{p}\in\ell^{t}(\mathbb{N})$ for any $t>p$, the use of this map yields the bounds

### v3

@@ 3 $m$-widths of $(\text{{{b}}},1)$-holomorphic functions: the results of [6]
Let $\mathcal{K}$ be a set of (holomorphic) functions. To understand how well we can approximate functions in $\mathcal{K}$ from finitely-many samples, we use the theory of $m$-widths (see, e.g., [31]). Recall that $\mathcal{V}$ is a Hilbert space, $\varrho$ is a tensor-product probability measure on $\mathcal{U}$ and consider the Lebesgue–Bochner space $\mathcal{X}=L^{2}_{\varrho}(\mathcal{U};\mathcal{V})$. Let $\mathcal{Y}=C(\mathcal{U};\mathcal{V})$ be the normed vector space of continuous functions from $\mathcal{U}$ to $\mathcal{V}$ with respect to the uniform norm. Then, following [6], we define the (adaptive) $m$-width
@@
Now consider $(\bm{b},\varepsilon)$-holomorphic functions. In [6], we made the distinction between the known anisotropy (known $\bm{b}$) and unknown anisotropy (unknown $\bm{b}$) cases. In the known anisotropy case, we now let $\mathcal{K}=\mathcal{H}(\bm{b})$ in (3.1) for some fixed $\bm{b}$, so that the reconstruction map $\mathcal{T}$ can depend on $\bm{b}$. Conversely, in the case of unknown anisotropy we let $\mathcal{K}=\mathcal{H}(p)$ or $\mathcal{K}=\mathcal{H}(p,\mathsf{M})$ in (3.1), in which case the reconstruction map is prohibited from depending on $\bm{b}$.
@@
In the unknown anisotropy case, we write
@@
We now recap the main results of [6]. These results only considered the case where the probability measure $\varrho=\varrho_{1}\times\varrho_{1}\times\cdots$ is a tensor-product of a single probability measure on $[-1,1]$ (the existence of this measure is guaranteed by the Kolmogorov extension theorem [23, Thm. 14.36]). However, they can be easily generalized to tensor-products of different one-dimensional measures, subject to the following assumption (see Remark 3).
@@
    The probability measure $\varrho$ on $\mathcal{U}$ has the form $\varrho=\varrho_{1}\times\varrho_{2}\times\cdots$, where each $\varrho_{i}$ is a probability measure on $[-1,1]$. Moreover, $\inf_{i}\sigma_{i}>0$, where the $\sigma_{i}$ are given by
    $\sigma_{i}=\sqrt{\int^{1}_{-1}(y-\upsilon_{i})^{2}\,\mathrm{d}\varrho_{i}(y)},\quad\text{for }\upsilon_{i}=\int^{1}_{-1}y\,\mathrm{d}\varrho_{i}(y).$
@@
The bound (3.5) pertains to the known anisotropy case. It shows that no sampling-recovery pair can achieve a rate better than $m^{1/2-1/p}$ for $\bm{b}\in\ell^{p}(\mathbb{N})$ or $\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N})$ when the reconstruction map is allowed to depend on $\bm{b}$. For the unknown anisotropy case, (3.6) shows that it is impossible to approximate functions in $\mathcal{H}(p)$ from finitely-many samples. This arises because functions in $\mathcal{H}(p)$ can have important variables in infinitely-many different configurations [6]. The final result (3.7) asserts that when restricting to the class $\mathcal{H}(p,\mathsf{M})$ one can achieve a rate of at best $m^{1/2-1/p}$.
We remark that [6, Thms. 4.6 & 4.7] also establish upper bounds for several of the $m$-widths considered above when $\varrho$ is the uniform measure. However, these bounds do not consider i.i.d. pointwise samples and are therefore not relevant to this work. In the next section, we present counterpart upper bounds based on such samples. In particular, we show that the rates $m^{1/2-1/p}$ can be ‘nearly’ achieved from i.i.d. pointwise samples drawn from general tensor-product Jacobi measures, using suitable recovery maps based on the corresponding tensor Jacobi polynomials.
@@
    The extension of the results of [6] to the case where $\varrho$ satisfies Assumption 3 involves a straightforward modification of several proofs. First, one defines the orthonormal (in $L^{2}_{\varrho}(\mathcal{U})$) functions $\psi_{i}(\bm{y})=\frac{y_{i}-\upsilon_{i}}{\sigma_{i}}$ and then observes that [6, Lem. 5.2] remains valid with $\sigma=\inf_{i}\sigma_{i}$ and $\tau=\sup_{i}|\upsilon_{i}|$. Observe that $|\tau|\leq 1$ since each $\varrho_{i}$ is a probability measure supported on $[-1,1]$. Lemmas 5.3 and 5.4 of [6] then extend without further modifications and with these values of $\sigma$ and $\tau$. Finally, the same applies to Theorems 4.4 and 4.5 of [6], on which Theorem 3.2 is based.
@@
    In general, $\overline{\theta_{m}}(p)\neq\overline{\theta_{m}}(p,\mathsf{M})$. However, $\overline{\theta_{m}}(p)=\overline{\theta_{m}}(p,\mathsf{M})$ whenever the measure $\varrho=\varrho_{1}\times\varrho_{1}\times\cdots$ is a tensor-product of a single measure $\varrho_{1}$ on $[-1,1]$. This follows from two observations. First, for such a measure, we have $\theta_{m}(\bm{b})=\theta_{m}(\pi(\bm{b}))$ for any permutation $\pi$. Second, if $\bm{b}\in\ell^{p}(\mathbb{N})$ with ${\|\bm{b}\|}_{p}\leq 1$, then there is a permutation $\pi$ for which $\pi(\bm{b})$ is nonincreasing, and therefore $\pi(\bm{b})\in\ell^{p}_{\mathsf{M}}(\mathbb{N})$ with ${\|\pi(\bm{b})\|}_{p,\mathsf{M}}={\|\bm{b}\|}_{p}$. See Appendix B for the proof of this claim, and the exact relationship between $\theta_{m}(\bm{b})$ and $\theta_{m}(\pi(\bm{b}))$ in the general case.
@@ 4.1 Jacobi measures
For $\alpha,\beta>-1$ the Jacobi measure on $[-1,1]$ is given by
@@
It is a short argument to show that Assumption 4.1 implies Assumption 3 for the measure (4.3). This follows immediately, after noticing that $\sigma_{i}$ is a continuous function of $\alpha_{i},\beta_{i}$ in this case. Therefore, the lower bounds of Theorem 3.2 apply in this setting.
@@ 4.2 Upper bounds in expectation
Let $\bm{\alpha},\bm{\beta}$ satisfy Assumption 4.1 with constant $\tau>0$ and $\varrho=\varrho_{\bm{\alpha},\bm{\beta}}$ be the tensor-product Jacobi probability measure as in (4.3). Let $m\geq 3$ and consider sample points $\bm{y}_{1},\ldots,\bm{y}_{m}\sim_{\mathrm{i.i.d.}}\varrho$. Then there is a reconstruction map $\mathcal{T}:(\mathcal{U}\times\mathcal{V})^{m}\rightarrow L^{2}_{\varrho}(\mathcal{U};\mathcal{V})$ depending on $\bm{\alpha}$ and $\bm{\beta}$ only such that
$\theta_{m}(\bm{b})\leq\mathbb{E}\left(\sup_{f\in\mathcal{H}(\bm{b})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\right)\leq C(\bm{b},p,\tau)\cdot\left(\frac{m}{\log^{5}(m)}\right)^{\frac{1}{2}-\frac{1}{p}},$
@@
$\theta_{m}(p,\mathsf{M})\leq\mathbb{E}\left(\sup_{f\in\mathcal{H}(p,\mathsf{M})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\right)\leq C(p,t,\tau)\cdot m^{\frac{1}{2}-\frac{1}{t}},$
@@
Note that, in both Theorems 4.2 and 4.3, the leftmost inequalities (namely, $\theta_{m}(\bm{b})\leq\ldots$ and so forth) follow immediately from the definition of the relevant quantities, i.e., (3.2)–(3.4).
@@ 4.3 Discussion
We now discuss some important aspects of Theorems 4.2 and 4.3.
Theorem 3.2 provides lower bounds for the various $m$-widths, and establishes that the rate $m^{1/2-1/p}$ is the best possible. These lower bounds are valid for Banach-valued functions, whereas our main results hold only when $\mathcal{V}$ is a Hilbert space. This gap was previously encountered in [4]. There, (nonuniform) upper bounds for Legendre and Chebyshev polynomial approximations were shown for Banach-valued functions. However, the algebraic rates of convergence shown therein are slower than those of the Hilbert-valued case. A similar extension of Theorems 4.2 and 4.3 to Banach spaces would also result in similar reductions of the rates. As of now, it is unknown how to attain upper bounds for the $m$-width for Banach-valued functions that match the lower bounds.
In the second parts of Theorems 4.2 and 4.3, we trade the (poly)logarithmic factor from the first part for an (arbitrarily small) reduction in the algebraic rate in order to obtain an upper bound for the corresponding $m$-width. This is done because it is not possible, with our current proof techniques, to bound the constant $C(\bm{b}_{\pi},p,\tau)$ (respectively, $C(\bm{b},p,\tau)$) uniformly over sequences $\bm{b}\in\ell^{p}(\mathbb{N})$ with ${\|\bm{b}\|}_{p}\leq 1$ (respectively, $\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N})$ with ${\|\bm{b}\|}_{p,\mathsf{M}}\leq 1$). Instead, we use an idea from [6] which allows us to bound $\sup_{{\|\bm{b}\|}_{p}\leq 1}C(\bm{b}_{\pi},t,\tau)<\infty$ for $t\in(p,1)$ and likewise for $C(\bm{b},t,\tau)$. See the proofs of Theorems 4.2 and 4.3 in §6 for further details.
@@
As noted previously, the unknown anisotropy case studied in Theorem 4.3 requires a different approach to the least-squares approximation used in Theorem 4.2. Due to the intrinsic unknown nature of the entries of $\bm{b}$, we use techniques based on compressed sensing to establish theoretical guarantees on the recovery of certain polynomial coefficients in a large (but finite) index set. A key step involves using well-established results that assert the (weighted) Restricted Isometry Property (RIP) for the relevant measurement matrix. The need to assert the RIP leads to a $\log^{4}(m)$ factor in the first bound in Theorem 4.3. Unfortunately, it is well-known that improving the log factors in RIP-type estimates is a challenging open problem. See [10] for further discussion.
  • Remark 4.4
    The one additional power of $\log(m)$ in Theorem 4.3, which results in $\log^{4}(m)\cdot\log(m)=\log^{5}(m)$, stems from the desire to have a reconstruction map $\mathcal{T}$ that is valid for all $10<p<$ 1. In fact, the term $\log^{5}(m)$ can be replaced by $\log^{4}(m)g(m)$, where $g$ is any increasing function $g:\mathbb{N}\rightarrow[1,\infty)$ with $g(m)\rightarrow+\infty$ as $m\rightarrow+\infty$ and which grows subalgebraically fast in $m$. We use $g(m)=\log(m)$ to simplify the presentation. In the general case, the constants $C$ also depend on $g$. See §6.3 and §7.
While Theorem 4.3 relies on compressed sensing, it is also possible to derive a weaker version of it by using the least-squares approximation from Theorem 4.2. Notice that any $\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N})$, $\bm{b}\geq\bm{0}$, ${\|\bm{b}\|}_{p,\mathsf{M}}\leq 1$, satisfies $\bm{b}\leq\bm{b}_{p}$, where $\bm{b}_{p}=(i^{-1/p})_{i\in\mathbb{N}}$. It follows from (2.3) that $\mathcal{R}(\bm{b}_{p},\varepsilon)\subseteq\mathcal{R}(\bm{b},\varepsilon)$, which implies that $\mathcal{H}(\bm{b})\subseteq\mathcal{H}(\bm{b}_{p})$ and, in turn, that $\theta_{m}(\bm{b})\leq\theta_{m}(\bm{b}_{p})$. Now consider the reconstruction map $\mathcal{T}=\mathcal{T}_{p}$ from Theorem 4.2 based on $\bm{b}_{p}$. Since $\bm{b}_{p}\in\ell^{t}(\mathbb{N})$ for any $t>p$, the use of this map yields the bounds